1. Pretend pretzels and chips are perfect substitutes. Billy has $10. If chips cost $4 and pretzels cost $3 what is Billy's demand for chips?

2. At what quantity will the marginal cost curve cross the average cost curve? C = 3q - 4q^2 + q^3.

3. Jessica's income elasticity of demand for donuts is e=-2. If Jessica gets a 3% cost of living adjustment, what happens to her demand for donuts?

a) The vertical intercept for demand curve moves down by 6%

b) The vertical intercept for demand curve moves down by 3%

c) The vertical intercept for demand curve moves up by 6%

d) The vertical intercept for demand curve moves up by 3%

e) Can’t be determined

4. Does the function Q=5LK^½ have diminishing marginal returns to labor in the SR?
Answer #1

1) it will be zero units because both are perfect substitutes which means consumer will prefer only the cheaper product. In this case the cheaper product is pretzels so only pretzels will be purchased

2) use MC = AC

3 - 8q + 3q^2 = 3 - 4q + q^2

2q^2 - 4q = 0

This implies q = 2.

3) option A is correct

4) no. It has increasing returns.